cosx+cosy-cos(x+y)=3/2

Grade:10

1 Answers

Arun Kumar IIT Delhi
9 years ago
Hello Student,

cosx+cosy-cos(x+y)=3/2
2cos((x+y)/2)cos((x-y)/2)-2[cos((x+y)/2))]^2-0.5=0
=>2ka-2k^2-0.5=0
=>D=4a^2-4>=0
=>a=1 since its a cos(x-y) whose value is in [-1,1]
so x=y
You got what you asked for.

Thanks & Regards
Arun Kumar
Btech, IIT Delhi
Askiitians Faculty